# On Meta-Prompting

### v2

@@
Abstract.
Modern generative language models are capable of interpreting input strings as instructions, or prompts, and carry out tasks based on them. Many approaches to prompting and pre-training these models involve the automated generation of these prompts: meta-prompting, or prompting to obtain prompts. We propose a theoretical framework based on category theory to generalize and describe them. This framework is flexible enough to account for stochasticity, and allows us to obtain formal results around task agnosticity and equivalence of various meta-prompting approaches. Experimentally, we test our framework in two active areas of model research: creativity and ideation. We find that user preference strongly favors ($p<0.01$) the prompts generated under meta-prompting, as well as their corresponding outputs, over a series of hardcoded baseline prompts that include the original task definition. Using our framework, we argue that meta-prompting is more effective than basic prompting at generating desirable outputs.
@@ Abstract.
ccs: Computing methodologies Natural language processingccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language generationccs: Theory of computation Machine learning theory
1. Introduction
@@
Instruction-pretrained models (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022) such as large language models (LLMs) are able to interpret input strings (prompts) as instructions to carry out a task. This has led to their deployment in multiple product areas, such as writing assistance. However, it is well-known that these models have shortcomings, amongst which is their sensitivity to the prompt’s phrasing (Khatun and Brown, 2023; Xiong et al., 2024; Lu et al., 2022). This has led to a considerable body of work to determine optimal prompting techniques and wordings, which in turn has been shown to be automatable (Zhou et al., 2023; Shin et al., 2020), and effective when the model has some sort of guidance (Wei et al., 2022b; Zhang et al., 2023) or feedback (Richards, 2023).
@@
This work is largely applied, and, to our knowledge, there is not theoretical work characterizing LLM prompting, task adaptability, and user interaction, in spite of this technology being so widespread. This is perhaps because, from a mathematical perspective, complexities also arise when attempting to model LLM behavior, at least when accounting for their generalizability to arbitrary tasks, their stochasticity, and the intrinsic black-box nature of very large neural networks.
Factoring in user interaction makes theoretical work especially difficult, given its fine-grained and personal nature. For example, it would be unrealistic to assume that a set of users desiring to do the same thing (e.g., paraphrasing their own document) will use the same prompt when interacting with the LLM. Nonetheless, all users will expect analogous experiences and performances.
In this work we address this gap by modeling prompting approaches and task-specific behavior of a fixed LLM by using category theory. We use this mathematical framework to generalize and characterize various automated prompting approaches under the meta-prompting, or prompting to obtain prompts, umbrella. Finally, we test out this framework’s predictions on realistic applications related to writing assistance, and show that they are consistent with what it is observed empirically.
1.1. Contributions
We have two core contributions in this work.
  1. (1)
    We provide a mathematical framework for LLMs that allows for the study of the relationships between prompting, task execution, and user interaction. With it, we show that meta-prompting approaches are task-agnostic processes that model language and user interaction without explicitly depending on the task. We also show that meta-prompting approaches are equivalent in a categorical sense. We also argue that this is a prompting technique that will always outperform traditional prompting.
  2. (2)
@@ 1.1. Contributions
    To provide evidence on this assertion, we experiment with meta-prompting by calling a basic meta-prompt over two common applications of LLMs related to writing assistance. We find that the prompts generated by meta-prompting are regarded by users as more suitable than baselines (hardcoded prompts and the original task description; $p<0.01$ under a Wilcoxon signed-rank test); and that, in line with our framework’s predictions, the outputs obtained by these generated prompts are likewise ranked as more suitable ($p<0.01$).
The first contribution is theoretical in nature, but its implications are of particular interest in agentic scenarios--that is, horizontal components (e.g., a chatbot) that interact with vertical components (e.g., specialists, such as a summarization component), all while using the same model. The second acts as an empirical test for our framework under relevant applications. Via our contributions we formally conclude that meta-prompting is generally better than fixed prompts at generating desirable outputs.
2. Background
2.1. Why Category Theory?
@@
Many aspects of ML, and LLMs in particular, rely on solid, well-understood foundations (for example, gradient descent), but it is not entirely clear how these foundations come together (Shiebler et al., 2021). For LLMs in particular, their size, sensitivity to the input, sometimes opaque design, and inconsistent (random) outputs [^1] Randomness in the output may be desirable for more diverse content, and hence better user experience. We do not consider it a flaw. means that mathematical modelling of these models is difficult. Additionally, a single LLM may have multiple distinct downstream applications with carefully-tuned task-specific prompts. Hence, developing the language required to model LLM behavior must be general enough to factor in their complexity, but also sufficiently expressive to account for their adaptability.
Category theory (Eilenberg and MacLane, 1945) is a branch of mathematics that aims to study analogies between different concepts in mathematics by abstracting out certain details, but retaining rigor. [^2] An apt analogy is that of Spivak (2014) (p. 376), who compares it to studying the game of tennis without studying the physics of every particle in the system. This makes it suitable as a language to describe LLMs.
2.2. Category Theory
@@ 2.2.1. Fundamentals
Although objects and structures in a category do not need to be sets--hence our distinction between hom objects and hom sets--in this paper we will only deal with locally small categories.
@@
2.2.2. Diagrams, Functors, and Natural Transformations
@@
Functors can be thought of as analogies between categories (Baez, 2006; Riehl, 2016).
More importantly, they can also be transformed between one another through a natural transformation $\alpha\colon F\implies G$. It is normally said that natural transformations are where the real power of category theory lies (Mac Lane, 1998).
@@
2.2.3. Structure and Classification of Categories
@@
The object $\text{hom}(X,Z)$ is an object in C, known as the exponential object if $\otimes$ is the Cartesian product, and written as $Z^{X}$. The notation $\text{{C}}^{\text{op}}$ denotes the same category as C, but with the arrows reversed--often known as the opposite category. If C is small, the exponential object is the set of all functions from $X$ to $Z$.
@@ 2.3. Further Assumptions
2.4. LLMs and Prompting
For our purposes, a LLM is a ‘‘box’’ that takes in strings, and according to some internal logic, outputs strings. [^3] Our reluctance to call them “functions” will be clear in this section: our morphisms are not the LLM itself, but the prompts. As mentioned earlier, in category theory not all morphisms are functions. The input strings can be phrased to be instructions, or prompts, and the model will execute them and return the desired output.
2.4.1. User and System Prompts
Contemporary LLMs separate the concepts of system and user prompts. The system prompt is usually developer-specified and designed to ensure responsible and reliable interaction, as well as to provide a task definition. The user prompt, on the other hand, is the user’s instructions along with the user-provided content.
For example, suppose a user types into a writing assistant ‘‘give me the gist of the following document, but do not use the letter ‘e’’’. In this case:
@@ 2.4.1. User and System Prompts
  • The user prompt is the string ‘‘give me the gist of the following document, but do not use the letter ‘e’’’.
  • The system prompt is a string such as ‘‘Summarize the document based on the user’s asks. Do not use any information that is not in the document. Document: {document}. User’s asks: {user prompt}. Return the output in JSON format’’.
  • The output is a summary of the text, parseable via JSON, and, ideally, with no hallucinations or using the letter ‘‘e’’.
While it seem be unrealistic at a first glance, the example above illustrates the distinctions between task definitions and user-provided constraints, along with the typical uniqueness and unpredictability of users in production systems.
We refer to the concatenation of the system and user prompt as the prompt, since the model observes both.
2.4.2. Assumptions on Prompt Execution
At runtime, the prompt sent to the model is the concatenation (or replacement) of the system and user prompts and inputs. We assume that the system prompt does not ever change. As we will see in the next sections, this is a weak assumption since the system prompt depends on the choice of category.
Contemporary autoregressive LLMs are stochastic (unless parametrized otherwise) and operate over tokens, not characters. They generate their output by mapping probabilities into tokens (decoding) starting from some token (e.g., $\langle s\rangle$), and stop when a pre-specified maximum output length, or another token denoting end of generation (e.g., $\langle/s\rangle$), has been reached.
@@ 2.4.2. Assumptions on Prompt Execution
We assume the LLM to be deterministic, that it has followed the instructions correctly, and that its output is grammatical. When the user input is not conducive to do so (e.g., ‘‘multiply {X} by $3$’’, where X is an extract of Finnegans Wake), we assume the LLM responds with a boilerplate string (e.g., ‘‘As a LLM I cannot perform this task’’). All of these assumptions are only for simplicity: we can always consider the codomain the entire $\Sigma^{k}$, or map any failed outputs to a special token.
2.4.3. A Mathematical Model of Prompting and Prompt-Sensitivity
@@
Two prompts $p,q$ describing the same task (e.g., $p:=\;$‘‘multiply {X} by 3’’ and $q:=\;$‘‘add {X} 3 times’’) do not have guaranteed equal behavior, as LLMs are prompt-sensitive. We consider isomorphism to be over meaning (i.e., two paraphrases without loss of meaning are isomorphic). Hence $p$ and $q$ have distinct codomains and so $\tilde{p}\neq\tilde{q}$. Prompt application is associative up to isomorphism: suppose $p:=``\text{Expand }\{X\}$’’, $q:=``\text{Get keypoints }\{Y\}$’’, and $r:=``\text{Translate to West Frisian }\{Z\}$’’. Then $(pq)r\cong p(qr)\cong pqr$.
3. A Categorical View of Prompting an LLM
3.1. The Prompt Category
@@
Prompt is a right-closed [^4] We chose right-closed for convenience, but symmetric results may be obtained if Prompt is left-closed. monoidal category with all subsets of $\Sigma^{k}$ as objects, and all possible instructions expressible as members of $\Sigma^{k}$ as arrows. Identity arrows are prompts of the form $1_{X}:=\;$‘‘Return {X}’’. Isomorphism is given over meaning. The composition operation is sequential prompt application.
@@
3.2. Task-Categories
While Prompt presents a generalized application of LLMs, in practice it is more desirable to model specific tasks, such as summarization or chatting. This is because in downstream applications a feature may have a carefully-tuned, task-specific system prompt (e.g., ‘‘you are a summarizer bot’’). Alternatively, because a general system prompt (e.g. ‘‘you are a helpful AI assistant’’) needs to comply with the user’s requests. This runtime execution may be phrased as a special kind of category, which we call a task-category.
@@ 3.2. Task-Categories
Note that we do not assume that Task is a full subcategory of Prompt--i.e., we do not assume $\text{Hom}_{\text{{Task}}}(X,Y)=\text{Hom}_{\text{{Prompt}}}(X,Y)$ for all $X,Y\in\text{{Task}}$.
@@
3.3. An Example
It may be helpful to illustrate this with an example. Let Summ be a task-category whose objects are passages; and the arrows are prompts that take passages to their summaries, such as $f(X):=\;$‘‘Summarize {X}’’. Since it is possible to obtain a summary of a summary, the arrows are composable and associative, and hence Summ is a category. We assign the internal hom to be the string descriptions of the prompts.
Let Expand be the task-category with key points for writing passages, and passages, as objects. Its arrows are prompts that take key points to passages, such as $f(X):=\;$‘‘Write a passage using the following: $\{X\}$’’. Composition and associativity are symmetric to Summ: all expanded key points are new key points.
@@ 3.3. An Example
This functor is easy to construct: consider a replacement of the form ‘‘instead of $\{\tilde{f}\}$, do $\{\tilde{f}^{*}\}$’’. We can build an analogous $G\colon\text{{Expand}}\rightarrow\text{{Summ}}$. A (monoidal) functor between two monoidal categories must preserve the monoidal structure, but we assume this implicitly since the tensor product and natural isomorphisms are the same for task-categories.
3.4. Task-Category Equivalence
@@
To do this, we introduce in Lemma 1 our first technical lemma, which formally shows that if it is possible to rephrase one task definition (i.e., morphisms) into another, then their task-categories are equivalent.
Lemma 0 (Equivalence of Task-Categories).
@@
To see this, note that every $t\in\text{Img}(f_{1})$ is a rewrite of $\tilde{T}_{1}$. Then every $h_{1}\in\text{Hom}_{\text{{Task}}_{1}}(-,-)$ as defined by $T_{1}$ is in a one-to-one correspondence with $\text{Img}(f_{1})$, along with a corresponding exponential object in $\text{{Task}}_{1}$ with the verbatim string description. A symmetric argument applies to $\text{{Task}}_{2}$. Hence the functor $F$ may be constructed by mapping back the exponential objects to every $h_{1}$ (r. $h_{2}$). ∎
Informally, the proof of Lemma 1 has two insights: first, that Rewrite represents both a natural language task and is also a tool for the formal analysis of task-categories. Second, that, if every prompt from $T_{1}$ can be rephrased into some $T_{2}$ (i.e., if there exists a natural transformation $\alpha\colon T_{1}\implies T_{2}$), then we may rephrase $\text{{Task}}_{1}$ into $\text{{Task}}_{2}$.
@@
4. Meta-Prompting
4.1. Meta-Prompting in Category Theory
The isomorphisms associated with the internal hom of a category are a map between morphisms of $\text{Hom}_{\text{{Task}}}(Y,Z^{X})$ and $\text{Hom}_{\text{{Task}}}(X\otimes Y,Z)$. Let $\lambda\colon Y\rightarrow Z^{X}$ be one such morphism. In a task-category, $Z^{X}$ is precisely the set of prompts from the input $X$ to the LLM output $Z$. When $X$ is the system prompt (task description) and $Y$ is the user-provided content, $\lambda$ selects any element from the prompt set $Z^{X}$ based on the user-provided content $Y$. From an applied perspective, this is a ‘‘box’’ (a prompt) that takes in contexts and return prompts: a meta-prompt morphism.
4.2. Properties of Meta-Prompting
4.2.1. Agnosticity
For a given task, meta-prompt morphisms are system prompt-agnostic. This is easy to see by noting that by construction all morphisms within the task-category are relevant to the task. Hence the meta-prompt morphism will always return an appropriate map between $X\rightarrow Z$ given $Y$. In other words, meta-prompt morphisms encode what the LLM should ‘‘do’’ given different descriptions of the same task.

### v3

@@
Abstract
Modern large language models (LLMs) are capable of interpreting input strings as instructions, or prompts, and carry out tasks based on them. Unlike traditional learners, LLMs cannot use back-propagation to obtain feedback, and condition their output in situ in a phenomenon known as in-context learning (ICL). Many approaches to prompting and pre-training these models involve the automated generation of these prompts, also known as meta-prompting, or prompting to obtain prompts. However, they do not formally describe the properties and behavior of the LLMs themselves. We propose a theoretical framework based on category theory to generalize and describe ICL and LLM behavior when interacting with users. Our framework allows us to obtain formal results around task agnosticity and equivalence of various meta-prompting approaches. Using our framework and experimental results we argue that meta-prompting is more effective than basic prompting at generating desirable outputs.
keywords:
@@ Abstract
1 Introduction
@@
Instruction-pretrained models (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022) such as large language models (LLMs) are able to interpret input strings (prompts) as instructions to carry out a task. Given the lack of a feedback cycle (e.g., back-propagation), this is known as in-context learning (ICL; Wei et al. 2022b; Okawa et al. 2023; Ganguli et al. 2022; Brown et al. 2020; to name a few). ICL has led their deployment in multiple product areas, such as writing assistance, in spite of their sensitivity to the prompt’s phrasing (Khatun and Brown, 2023; Xiong et al., 2024; Lu et al., 2022). In turn, this spawned a considerable body of work to determine optimal prompting techniques and wordings, which itself has been shown to be automatable (Zhou et al., 2023; Shin et al., 2020), and effective when the model has some sort of guidance (Wei et al., 2022c; Zhang et al., 2023; Richards, 2023).
@@
This work is largely applied, and, to our knowledge, there is not theoretical work characterizing LLM prompting, task adaptability, and user interaction in the context of ICL. This is perhaps because, from a mathematical perspective, complexities also arise when attempting to model LLM behavior, at least when accounting for their generalizability to arbitrary tasks, their stochasticity, and the intrinsic black-box nature of very large neural networks. Indeed, factoring in user interaction makes theoretical work especially difficult, given its fine-grained and personal nature. For example, it would be unrealistic to assume that a set of users desiring to do the same thing (e.g., paraphrasing their own document) will use the same prompt when interacting with the LLM. Nonetheless, all users will expect analogous experiences and performances.
In this work we address this gap by modeling ICL-based prompting approaches and task-specific behavior of a fixed LLM by using category theory. This framework allows for the study of the relationships between prompting, task execution as in-context learning, and user interaction, while factoring out issues such as prompt sensitivity and generalizability. With it, we show that meta-prompting approaches are task-agnostic processes that model language and user interaction. We show that meta-prompting approaches are equivalent in a categorical sense, and argue that they will always outperform traditional prompting. Our contribution is theoretical in nature, but its implications are of particular interest in agentic scenarios–that is, horizontal components (e.g., a chatbot) that interact with vertical components (e.g., specialists, such as a summarization component), all while using the same model.
@@
We provide a small set of experiments to provide evidence on this assertion in Appendix D. There we show that the prompts generated by meta-prompting are regarded by users as more suitable than baselines (hardcoded prompts and the original task description; $p<0.01$ under a Wilcoxon signed-rank test); and that, in line with our framework’s predictions, the outputs obtained by these generated prompts are likewise ranked as more suitable ($p<0.01$).
2 Background
2.1 Why Category Theory?
@@
Many aspects of ML, and LLMs in particular, rely on solid, well-understood foundations (for example, gradient descent), but it is not entirely clear how these foundations come together (Shiebler et al., 2021). For LLMs in particular, their size, sensitivity to the input, sometimes opaque design, and inconsistent (random) outputs [^1] Randomness in the output may be desirable for more diverse content, and hence better user experience. We do not consider it a flaw. means that mathematical modelling of these models is difficult. Additionally, a single LLM may have multiple distinct downstream applications with carefully-tuned task-specific prompts. Hence, developing the language required to model LLM behavior must be general enough to factor in their complexity, but also sufficiently expressive to account for their adaptability. Category theory (Eilenberg and MacLane, 1945) is a branch of mathematics that aims to study analogies between different concepts in mathematics by abstracting out certain details, but retaining rigor. [^2] An apt analogy is that of Spivak (2014) (p. 376), who compares it to studying the game of tennis without studying the physics of every particle in the system. This makes it suitable as a language to describe LLMs.
2.2 Category Theory
@@ 2.2.1 Fundamentals
Although objects and structures in a category do not need to be sets–hence our distinction between hom objects and hom sets–in this paper we will only deal with locally small categories.
@@
2.2.2 Diagrams, Functors, and Natural Transformations
@@
Functors can be thought of as analogies between categories (Baez, 2006; Riehl, 2016). More importantly, they can also be transformed between one another through a natural transformation $\alpha\colon F\implies G$. It is normally said that natural transformations are where the real power of category theory lies (Mac Lane, 1998).
@@
2.2.3 Structure and Classification of Categories
@@
The object $\text{hom}(X,Z)$ is an object in C, known as the exponential object if $\otimes$ is the Cartesian product, and written as $Z^{X}$. The notation $\text{{C}}^{\text{op}}$ denotes the same category as C, but with the arrows reversed–often known as the opposite category. If C is small, the exponential object is the set of all functions from $X$ to $Z$.
@@ 2.3 Further Assumptions
2.4 LLMs and Prompting
For our purposes, a LLM is a “box” that takes in strings, and according to some internal logic, outputs strings. [^3] Our reluctance to call them “functions” will be clear in this section: our morphisms are not the LLM itself, but the prompts. As mentioned earlier, in category theory not all morphisms are functions. The input strings can be phrased to be instructions, or prompts, and the model will execute them and return the desired output. In the context of ICL, an LLM can be viewed as a function $LLM\colon\Sigma\rightarrow\Sigma$, for a finite set of strings (tokens) $\Sigma$. In more detail, contemporary LLMs are autoregressive–that is, their next output $w_{k}$ is conditioned on the previously-generated tokens $w_{k-1},\dots w_{0}$, and the given input tokens $v_{0},\dots,v_{m}$ for some $k,m$, for tokens in $\Sigma$. Then their output can be described [^4] In practice, this is more complicated due to optimizations performed to calculate the output. as
$\text{Pr}[w_{k}]=\prod_{j}^{k-1}\text{Pr}[w_{j}|w_{i<j},\dots v_{m},\dots,v_{0}]$ (1)
From this perspective, formally describing the interactions between the user, the LLM, and ICL becomes difficult; particularly since it does not capture specificity of the desired use within ICL.
2.4.1 User and System Prompts
The aforementioned difficulty increases given that contemporary LLMs separate the concepts of system and user prompts, both described above as the sequence $v_{m},\dots,v_{0}$. The system prompt is usually developer-specified and designed to ensure responsible and reliable interaction, as well as to provide a task definition. The user prompt, on the other hand, is the user’s instructions along with the user-provided content.
For example, suppose a user types into a writing assistant “give me the gist of the following document, but do not use the letter ‘e”’. In this case:
@@ 2.4.1 User and System Prompts
  • The user prompt is the string “give me the gist of the following document, but do not use the letter ‘e”’.
  • The system prompt is a string such as “Summarize the document based on the user’s asks. Do not use any information that is not in the document. Document: {document}. User’s asks: {user prompt}. Return the output in JSON format”.
  • The output is a summary of the text, parseable via JSON, and, ideally, with no hallucinations or using the letter “e”.
While it seem be unrealistic at a first glance, the example above illustrates the distinctions between task definitions and user-provided constraints, along with the typical uniqueness and unpredictability of users in production systems. We refer to the concatenation of the system and user prompt as the prompt, since the model observes both.
2.4.2 Assumptions on Prompt Execution
At runtime, the prompt sent to the model is the concatenation (or replacement) of the system and user prompts and inputs. We assume that the system prompt does not ever change. As we will see in the next sections, this is a weak assumption since the system prompt depends on the choice of category. As seen earlier, contemporary autoregressive LLMs are stochastic (unless parametrized otherwise) and operate over tokens, not characters. They generate their output by mapping probabilities into tokens (decoding) starting from some token (e.g., $\langle s\rangle$), and stop when a pre-specified maximum output length, or another token denoting end of generation (e.g., $\langle/s\rangle$), has been reached.
@@ 2.4.2 Assumptions on Prompt Execution
We assume the LLM to be deterministic, that it has followed the instructions correctly, and that its output is grammatical. When the user input is not conducive to do so (e.g., “multiply {X} by $3$”, where X is an extract of Finnegans Wake), we assume the LLM responds with a boilerplate string (e.g., “As a LLM I cannot perform this task”). All of these assumptions are only for simplicity: we can always consider the codomain the entire $\Sigma^{k}$, or map any failed outputs to a special token.
2.4.3 A Mathematical Model of Prompting and Prompt-Sensitivity
@@
Two prompts $p,q$ describing the same task (e.g., $p:=\;$“multiply {X} by 3” and $q:=\;$“add {X} 3 times”) do not have guaranteed equal behavior, as LLMs are prompt-sensitive. We consider isomorphism to be over meaning (i.e., two paraphrases without loss of meaning are isomorphic). Hence $p$ and $q$ have distinct codomains and so $\tilde{p}\neq\tilde{q}$. Prompt application is associative up to isomorphism: suppose $p:=``\text{Expand }\{X\}$”, $q:=``\text{Get keypoints }\{Y\}$”, and $r:=$“Translate to West Frisian $\{Z\}$”. Then $(pq)r\cong p(qr)\cong pqr$.
3 A Categorical View of Prompting an LLM
3.1 The Prompt Category
@@
Prompt is a right-closed [^5] We chose right-closed for convenience, but symmetric results may be obtained if Prompt is left-closed. monoidal category with all subsets of $\Sigma^{k}$ as objects, and all possible instructions expressible as members of $\Sigma^{k}$ as arrows. Identity arrows are prompts of the form $1_{X}:=\;$“Return {X}”. Isomorphism is given over meaning. The composition operation is sequential prompt application.
@@
3.2 Task-Categories
While Prompt presents a generalized application of LLMs, in practice it is more desirable to model specific tasks, such as summarization or chatting. This is because in downstream applications a feature may have a carefully-tuned, task-specific system prompt (e.g., “you are a summarizer bot”). Alternatively, because a general system prompt (e.g. “you are a helpful AI assistant”) needs to comply with the user’s requests. This runtime execution may be phrased as a special kind of category, which we call a task-category.
@@ 3.2 Task-Categories
Note that we do not assume that Task is a full subcategory of Prompt–i.e., we do not assume $\text{Hom}_{\text{{Task}}}(X,Y)=\text{Hom}_{\text{{Prompt}}}(X,Y)$ for all $X,Y\in\text{{Task}}$.
@@
3.3 An Example
It may be helpful to illustrate this with an example. Let Summ be a task-category whose objects are passages; and the arrows are prompts that take passages to their summaries, such as $f(X):=\;$“Summarize {X}”. Since it is possible to obtain a summary of a summary, the arrows are composable and associative, and hence Summ is a category. We assign the internal hom to be the string descriptions of the prompts.
Let Expand be the task-category with key points for writing passages, and passages, as objects. Its arrows are prompts that take key points to passages, such as $f(X):=\;$“Write a passage using the following: $\{X\}$”. Composition and associativity are symmetric to Summ: all expanded key points are new key points.
@@ 3.3 An Example
This functor is easy to construct: consider a replacement of the form “instead of $\{\tilde{f}\}$, do $\{\tilde{f}^{*}\}$”. We can build an analogous $G\colon\text{{Expand}}\rightarrow\text{{Summ}}$. A (monoidal) functor between two monoidal categories must preserve the monoidal structure, but we assume this implicitly since the tensor product and natural isomorphisms are the same for task-categories.
3.4 Task-Category Equivalence
@@
To do this, we introduce in Lemma 3.1 our first technical lemma, which formally shows that if it is possible to rephrase one task definition (i.e., morphisms) into another, then their task-categories are equivalent.
Lemma 3.1 (Equivalence of Task-Categories).
@@
To see this, note that every $t\in\text{Img}(f_{1})$ is a rewrite of $\tilde{T}_{1}$. Then every $h_{1}\in\text{Hom}_{\text{{Task}}_{1}}(-,-)$ as defined by $T_{1}$ is in a one-to-one correspondence with $\text{Img}(f_{1})$, along with a corresponding exponential object in $\text{{Task}}_{1}$ with the verbatim string description. A symmetric argument applies to $\text{{Task}}_{2}$. Hence the functor $F$ may be constructed by mapping back the exponential objects to every $h_{1}$ (r. $h_{2}$).
Informally, the proof of Lemma 3.1 has two insights: first, that Rewrite represents both a natural language task and is also a tool for the formal analysis of task-categories. Second, that, if every prompt from $T_{1}$ can be rephrased into some $T_{2}$ (i.e., if there exists a natural transformation $\alpha\colon T_{1}\implies T_{2}$), then we may rephrase $\text{{Task}}_{1}$ into $\text{{Task}}_{2}$.
@@
4 Meta-Prompting
4.1 Meta-Prompting in Category Theory
The isomorphisms associated with the internal hom of a category are a map between morphisms of $\text{Hom}_{\text{{Task}}}(Y,Z^{X})$ and $\text{Hom}_{\text{{Task}}}(X\otimes Y,Z)$. Let $\lambda\colon Y\rightarrow Z^{X}$ be one such morphism. In a task-category, $Z^{X}$ is precisely the set of prompts from the input $X$ to the LLM output $Z$. When $X$ is the system prompt (task description) and $Y$ is the user-provided content, $\lambda$ selects any element from the prompt set $Z^{X}$ based on the user-provided content $Y$. From an applied perspective, this is a “box” (a prompt) that takes in contexts and return prompts: a meta-prompt morphism.
4.2 Properties of Meta-Prompting
4.2.1 Agnosticity
For a given task, meta-prompt morphisms are system prompt-agnostic. This is easy to see by noting that by construction all morphisms within the task-category are relevant to the task. Hence the meta-prompt morphism will always return an appropriate map between $X\rightarrow Z$ given $Y$. In other words, meta-prompt morphisms encode what the LLM should “do” given different descriptions of the same task.